# Hölder curves with exotic tangent spaces

### v2

@@ 6. Hölder parametrizations of carpets $K^{\eta}$
By the arguments in the proof of Lemma 4.3, for each $w\in\mathcal{A}_{n}^{*}$ we have that $\phi_{w}^{\eta}(\partial[0,1]^{2})\subset K^{\eta}$. Therefore,
@@
Suppose now that the claim holds for some $m\in\mathbb{N}$ and let $u\in V_{m+1}$ be a vertex. Then there is a unique word $w\in\mathcal{A}_{n}^{m}$ so that $u\in\phi_{w}^{\eta}(T^{\eta(w)})$, and $u$ is a leaf in $T_{m+1}$ if and only if $(\phi_{w}^{\eta})^{-1}(u)$ is a leaf in $T^{\eta(w)}$, which is a leaf if and only if it has no edges in $T^{\eta(w)}$ emanating to the right or upward. This holds if and only if $u$ has no edges emanating upward or to the right in $T_{m+1}$. ∎
@@
We proceed by induction on $m$. For $m=1$, the result follows by simple inspection, as either $T_{1}=T^{1}$ or $T_{1}=T^{2}$. Now assume the result holds for some $m\in\mathbb{N}$. Let $(p,q)\subset T_{m+1}$ be an edge. Then there exist words $u,v\in\mathcal{A}_{n}^{m+1}$ so that $p=\phi_{u}^{\eta}({\bf 0})$ and $q=\phi_{v}^{\eta}({\bf 0})$. If there is a word $w\in\mathcal{A}_{n}^{m}$ so that $(p,q)\subset\phi_{w}^{\eta}(T^{\eta(w)})$, then there is some edge $e\subset T^{\eta(w)}$ so that $(p,q)=\phi_{w}^{\eta}(e)$, and the result follows from the $m=1$ case and from the definition of $\phi_{w}^{\eta}$.
If there is no such word $w$, then we have that $(p,q)\subset T_{m}$. Since the only nondegenerate line segments contained in $T_{m}$ are contained in its edges, there is some edge $e\subset T_{m}$ so that $(p,q)\subset e$, and since $(p,q)$ is not a subset of $\phi_{v}^{\eta}(T^{\eta(v)})$ for any $v\in\mathcal{A}_{n}^{m}$, we have that $[p,q]$ is not a subset of $\phi_{v}^{\eta}([0,1)^{2})$ for any $v\in\mathcal{A}_{n}^{m}$. Then by the induction hypothesis there is a word $w\in\mathcal{A}^{m}_{n}$ so that $e=\phi_{w}^{\eta}(e_{b})$ or $e=\phi_{w}^{\eta}(e_{l})$. Note that the length of $(p,q)$ is $n^{-m-1}$, that $p,q]\subset\phi_{w}^{\eta}([0,1]^{2})$, and that $(p,q)\subset\phi_{w}^{\eta}([0,1)^{2})$. Then either $p$ or $q$ is $\phi_{w}^{\eta}((0,1))$ or $\phi_{w}^{\eta}((1,0))$ since the edge $e$ is equal to either $\phi_{w}^{\eta}(e_{b})$ or to $\phi_{w}^{\eta}(e_{l})$. Then there is a $j\in\mathcal{A}_{n}$ so that the other of $p$ and $q$ is equal to $\phi_{wj}^{\eta}({\bf 0})$, so $\phi_{wj}^{\eta}(e_{b})=(p,q)$ or $\phi_{wj}^{\eta}(e_{l})=(p,q)$. ∎
In the next lemma, which is the crux of the proof of Proposition 6.1, we construct intermediate parametrizations $f_{m}:[0,1]\to T_{m}$.
@@
    For each $m\geq 0$ and $J\in\mathscr{E}_{m}$, there exists $w\in\mathcal{A}_{n}^{m}$ such that $f_{m}|J$ is a linear bijection onto either $\phi^{\eta}_{w}(e_{b})$ or $\phi^{\eta}_{w}(e_{l})$, and there exists a unique interval $J^{\prime}\in\mathscr{E}_{m}\setminus\{J\}$, so that $f_{m}|J^{\prime}=(f_{m}|J)\circ\zeta_{J^{\prime}}$ where $\zeta_{J^{\prime}}:J^{\prime}\to J$ is the unique linear orientation reversing map. Furthermore, $f_{m+1}(J)\subset\phi^{\eta}_{w}([0,1)^{2})$. Conversely, if $(p,q)$ is an edge of $T_{m}$, then there exist exactly two intervals $J,J^{\prime}\in\mathscr{E}_{m}$ such that $f_{m}(J)=f_{m}(J^{\prime})=(p,q)$.
@@
    There are disjoint nondegenerate closed intervals $I_{i,1},I_{i,2},I_{i,3},I_{i,4}\subset[0,1]$, denoted by $I_{i,j}=[a_{i,j},b_{i,j}]$, $j=1,\dots,4$ equal to the preimages of the four leaves of $T^{i}$, so that
@@
  • If $\tau\in\{\tilde{\tau}_{1},\tilde{\tau}_{2}\}$, then for a vertex $v\in V$ that is not a leaf, choose a nondegenerate component $I_{v}$ of the preimage $\tau^{-1}(\{v\})$, and for each vertex $v\in V$ that is a leaf other than $(1,0)$ or $(0,1)$, let $I_{v}=I_{i,j}$ where $j\in\{1,2,3,4\}$ is chosen so that $\tau(I_{i,j})=v$. We then define $\mathcal{N}(\tau)=\{I_{v}:v\in V\setminus\{(1,0),(0,1)\}\}$.
@@
Intervals in $\mathscr{E}_{m}$. Let $I\in\mathscr{E}_{m}$ and write $I=(a,b)$. By (P3) there exists $I^{\prime}=(a^{\prime},b^{\prime})\in\mathscr{E}_{m}\setminus\{I\}$ such that $f_{m}|I^{\prime}=(f_{m}|I)\circ\zeta_{I^{\prime}}$ where $\zeta_{I^{\prime}}:I^{\prime}\to I$ is the unique linear orientation reversing map and there exists a word $v\in\mathcal{A}_{n}^{m}$ so that $f_{m}(I)=\phi_{v}^{\eta}(e_{b})$ or $f_{m}(I)=\phi_{v}^{\eta}(e_{l})$. If $f_{m}(I)=\phi^{\eta}_{v}(e_{l})$, then, without loss of generality, we may assume that the $y$-coordinate of $f_{m}$ is increasing inside $I$ and decreasing inside $I^{\prime}$. Let $\tilde{\theta}_{I}:I\to(0,s_{\eta(v),U})$ be linear, bijective, and increasing, and let $\tilde{\xi}_{I^{\prime}}:I^{\prime}\to(s_{\eta(v),U},\tilde{t}_{\eta(v)})$ be linear, bijective, and decreasing. We define
@@
If $J\in\mathscr{N}_{m+1}(I)$, then there is a character $j\in\mathcal{A}_{n}$ so that $f_{m+1}(J)=\phi_{vj}^{\eta}({\bf 0})$, and we define $\operatorname{{\bf w}}_{m+1}(I)=vj$. Note that if $J,Q\in\mathscr{N}_{m+1}(I)\cup\mathscr{N}_{m+1}(I^{\prime})$ are distinct, then $\operatorname{{\bf w}}_{m+1}(J)\neq\operatorname{{\bf w}}_{m+1}(Q)$ as $\tilde{\theta}_{I}(J)\neq\tilde{\theta}_{I}(Q)$ are distinct intervals in, for example, $\mathcal{N}(\tilde{\tau}_{\eta(v)})$ ($\tilde{\theta}_{I}$ may be $\tilde{\xi}_{I^{\prime}}$ instead).
@@
By definition of $\tau_{\eta(\operatorname{{\bf w}}_{m}(I))}$, for any $Q\in\mathscr{N}_{m+1}(I)$, there is a character $j\in\mathcal{A}_{n}$ so that $f_{m+1}(Q)=\{\phi^{\eta}_{\operatorname{{\bf w}}_{m}(I)j}({\bf 0})\}$, and we set $\operatorname{{\bf w}}_{m+1}(Q)=\operatorname{{\bf w}}_{m}(I)j$.
@@
To conclude the proof, we show that $\operatorname{{\bf w}}_{m+1}:\mathscr{N}_{m+1}\to\mathcal{A}_{n}^{m+1}$ is surjective. Let $u\in\mathcal{A}_{n}^{m+1}$ be a word. By (P1) at $m$, there exists a unique interval $I\in\mathscr{N}_{m}$ so that $\operatorname{{\bf w}}_{m}(I)=u(m)$. If $f_{m}(I)$ is a leaf in $T_{m}$, then by definition of $\mathcal{N}(\tau_{i})$, for each $j\in\mathcal{A}_{n}$ there is an interval $J\in\mathscr{N}_{m+1}(I)$ so that $\operatorname{{\bf w}}_{m+1}(J)=\operatorname{{\bf w}}_{m+1}(I)j$, in particular one of them has $\operatorname{{\bf w}}_{m+1}(J)=u$. If $f_{m}(I)$ is not a leaf in $T_{m}$, then the result follows by definitions of $\mathcal{N}(\tilde{\tau}_{i})$ and $\mathcal{N}(\tau_{i})$ as well as (P3) at $m$. ∎
@@
It is possible that $\mathscr{N}_{m+1}\cap\mathscr{N}_{m}\neq\emptyset$ for some $m\in\mathbb{N}$. Define a map $\operatorname{{\bf w}}:\bigcup_{m\geq 0}\mathscr{N}_{m}\to\mathcal{A}_{n}^{*}$ as follows. Set $\operatorname{{\bf w}}([0,1])=\varepsilon$. If $m\in\mathbb{N}$ and $I\in\mathscr{N}_{m}\setminus\mathscr{N}_{m-1}$, set $\operatorname{{\bf w}}(I)=\operatorname{{\bf w}}_{m}(I)$. If $m\in\mathbb{N}$ and $I\in\mathscr{N}_{m}\cap\mathscr{N}_{m-1}$, set $\operatorname{{\bf w}}(I)=\operatorname{{\bf w}}_{m-k}(I)$, where $k\in\mathbb{N}$ is the greatest integer so that $I\in\mathscr{N}_{m-k}$. Note that $\operatorname{{\bf w}}$ is injective and $V_{m}=\{\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0}):I\in\mathscr{N}_{m}\}$ for every $m\in\mathbb{N}$.
@@
Fix $m\in\mathbb{N}$ and $x\in[0,1]$. By (P1), there exists $I\in\mathscr{N}_{m}\cup\mathscr{E}_{m}\cup\mathscr{F}_{m}$ so that $x\in I$. If $I\in\mathscr{F}_{m}$, then by (P5) we have that $f_{m}(x)=f_{m+1}(x)$. If $I\in\mathscr{N}_{m}$, then by (P4) we have that $f_{m}(x),f_{m+1}(x)\in\phi^{\eta}_{\operatorname{{\bf w}}_{m}(I)}([0,1]^{2})$. Finally, if $I\in\mathscr{E}_{m}$, then by (P3) we have that there is some word $v\in\mathcal{A}_{n}^{m}$ such that $f_{m}(x),f_{m+1}(x)\in\phi^{\eta}_{v}([0,1]^{2})$. In any case, we have the claimed estimate.
Hence, the functions $f_{m}$ converge uniformly to a continuous map $f:[0,1]\to[0,1]^{2}$. Furthermore, because the dendrites $T_{m}\subset K^{\eta}$ are nested and $f_{m}:[0,1]\to T_{m}$ are surjective, we have that the uniform limit $f$ is a continuous parametrization of $K^{\eta}$, as desired. ∎
@@
To prove porosity, note first that by (P3), (P4), and (P6) there exists a word $v\in\mathcal{A}_{n}^{m}$ such that $f(I)\cap f(J)\subset K^{\eta}_{v}$. Fix $x\in K^{\eta}$ and $r\in(0,n^{-m}]$. There exists an integer $k_{0}\in\mathbb{N}$ such that $n^{-k_{0}}\leq r<n^{-k_{0}+1}$, and so there exists $w\in\mathcal{A}_{n}^{k_{0}+1}$ with $K^{\eta}_{w}\subset\overline{B}(x,r)\cap K^{\eta}$.
@@
Therefore, we may assume for the rest of Case 2 that $u=v$ and that neither of $f(I),f(J)$ is a singleton. If $f_{k_{0}+1}(J)\cap K^{\eta}_{w}=\emptyset$, then $f(J)\cap\phi^{\eta}_{w}((0,1)^{2})=\emptyset$ by Lemma 4.2 and by (P3), (P4), and (P5), so the existence of such characters $i,j\in\mathcal{A}_{n}$ is clear. Similarly if $f_{k_{0}+1}(I)\cap K^{\eta}_{w}=\emptyset$. Now assume that $f_{k_{0}+1}(I)\cap K^{\eta}_{w}\neq\emptyset$ and $f_{k_{0}+1}(J)\cap K^{\eta}_{w}\neq\emptyset$. Then there are distinct leaves $p_{1},p_{2}\in T_{k_{0}+2}$ with $p_{1},p_{2}\in K^{\eta}_{w}$ and distinct intervals $Q_{1},Q_{2}\in\mathscr{N}_{k_{0}+2}$ such that $Q_{1}\subset J$ and $Q_{2}\subset I$ satisfy $f_{k_{0}+2}|Q_{1}\equiv p_{1}$ and $f_{k_{0}+2}|Q_{2}\equiv p_{2}$. Then $f([0,1]\setminus Q_{1})\cap\phi^{\eta}_{\operatorname{{\bf w}}_{k_{0}+2}(Q_{1})}((0,1)^{2})=\emptyset$ and $\operatorname{{\bf w}}(Q_{1})(k_{0}+1)=w$, so there is a pair of characters $i,j\in\mathcal{A}_{n}$ so that $wj=\operatorname{{\bf w}}_{k_{0}+2}(Q_{1})$ and $K^{\eta}_{wji}\subset\phi^{\eta}_{wj}((0,1)^{2})$. Since $I,J$ are distinct, and thus disjoint, we have that $f(I)\cap\phi^{\eta}_{wj}((0,1)^{2})=\emptyset$, concluding Case 2.
@@
Note that by (P7), there is some $I\in\mathscr{N}_{m+1}$ so that $f_{m+1}(I)$ is a leaf in $T_{m+1}$ and $I\subset J$. Then since $f_{m+1}(I)=\phi^{\eta}_{\operatorname{{\bf w}}(I)}({\bf 0})$ is a leaf in $T_{m+1}$, we have that
@@
If $Q\in\mathscr{F}_{m+1}$, then by (P5) there exists a word $u\in\mathcal{A}_{n}^{m+1}$ so that
@@
If $Q\in\mathscr{E}_{m+1}$, then by (P3) and Lemma 6.4, there exists a word $u\in\mathcal{A}_{n}^{m+1}$ so that either $f_{m+1}(Q)=\phi_{u}^{\eta}(e_{b})$ or $f_{m+1}(Q)=\phi_{u}^{\eta}(e_{l})$. In either case, by (P3), (P4), and (P5) we have that $f_{k}(Q)\subset\phi_{u}^{\eta}([0,1)^{2})$. Note that $\phi_{\operatorname{{\bf w}}(I)}^{\eta}(e_{l})$ emanates from $\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0})$ upward and $\phi_{\operatorname{{\bf w}}(I)}^{\eta}(e_{b})$ emanates from $\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0})$ to the right. Since $\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0})$ is a leaf in $T_{m+1}$, by Lemma 6.3 we must have that $u\neq\operatorname{{\bf w}}(I)$.
@@
so $\phi^{\eta}_{\operatorname{{\bf w}}(I)}((0,1)^{2})\cap K^{\eta}\subset f(I)$. Since $I$ is compact, we have $K^{\eta}_{\operatorname{{\bf w}}(I)}\subset f(I)$, so $K^{\eta}_{\operatorname{{\bf w}}(I)}\subset f(J)$ and $|\operatorname{{\bf w}}(I)|=m+1$. Therefore, $\pi_{\eta}\#\nu_{n}(f(J))\geq(5n-6)^{-m-1}$. ∎
@@
Let $f_{m}$, $\mathscr{E}_{m},\mathscr{N}_{m},\mathscr{F}_{m}$, and $\operatorname{{\bf w}}$ be as in Lemma 6.5 and Remark 6.6. There exists a non-decreasing surjection $\zeta:[0,1]\to[0,1]$ so that

### v3

@@ 6. Hölder parameterizations of carpets $K^{\eta}$
By Lemma 4.3, for each $w\in\mathcal{A}_{n}^{*}$ we have that $\phi_{w}^{\eta}(\partial[0,1]^{2})\subset K^{\eta}$. Therefore,
@@
Suppose now that the claim holds for some $m\in\mathbb{N}$ and let $u\in V_{m+1}$ be a vertex. Then there is a unique word $w\in\mathcal{A}_{n}^{m}$ such that $u\in\phi_{w}^{\eta}(T^{\eta(w)})$, and $u$ is a leaf in $T_{m+1}$ if and only if $(\phi_{w}^{\eta})^{-1}(u)$ is a leaf in $T^{\eta(w)}$, which is a leaf if and only if it has no edges in $T^{\eta(w)}$ emanating to the right or upward. This holds if and only if $u$ has no edges emanating upward or to the right in $T_{m+1}$. ∎
@@
We proceed by induction on $m$. For $m=1$, the result follows by simple inspection, as either $T_{1}=T^{1}$ or $T_{1}=T^{2}$. Now assume the result holds for some $m\in\mathbb{N}$. Let $(p,q)\subset T_{m+1}$ be an edge. Then there exist words $u,v\in\mathcal{A}_{n}^{m+1}$ satisfying $p=\phi_{u}^{\eta}({\bf 0})$ and $q=\phi_{v}^{\eta}({\bf 0})$. If there is a word $w\in\mathcal{A}_{n}^{m}$ such that $(p,q)\subset\phi_{w}^{\eta}(T^{\eta(w)})$, then there is some edge $e\subset T^{\eta(w)}$ with $(p,q)=\phi_{w}^{\eta}(e)$, and the result follows from the $m=1$ case and from the definition of $\phi_{w}^{\eta}$.
If there is no such word $w$, then we have that $(p,q)\subset T_{m}$. Since the only nondegenerate line segments contained in $T_{m}$ are contained in its edges, there is some edge $e\subset T_{m}$ such that $(p,q)\subset e$, and since $(p,q)$ is not a subset of $\phi_{v}^{\eta}(T^{\eta(v)})$ for any $v\in\mathcal{A}_{n}^{m}$, we have that $[p,q]$ is not a subset of $\phi_{v}^{\eta}([0,1)^{2})$ for any $v\in\mathcal{A}_{n}^{m}$. Then by the induction hypothesis there is a word $w\in\mathcal{A}^{m}_{n}$ satisfying either $e=\phi_{w}^{\eta}(e_{b})$ or $e=\phi_{w}^{\eta}(e_{l})$. Note that the length of $(p,q)$ is $n^{-m-1}$, that $p,q]\subset\phi_{w}^{\eta}([0,1]^{2})$, and that $(p,q)\subset\phi_{w}^{\eta}([0,1)^{2})$. Then either $p$ or $q$ is $\phi_{w}^{\eta}((0,1))$ or $\phi_{w}^{\eta}((1,0))$ since the edge $e$ is equal to either $\phi_{w}^{\eta}(e_{b})$ or to $\phi_{w}^{\eta}(e_{l})$. Then there is a $j\in\mathcal{A}_{n}$ such that the other of $p$ and $q$ is equal to $\phi_{wj}^{\eta}({\bf 0})$, therefore $\phi_{wj}^{\eta}(e_{b})=(p,q)$ or $\phi_{wj}^{\eta}(e_{l})=(p,q)$. ∎
In the next lemma, which is the crux of the proof of Proposition 6.1, we construct intermediate parameterizations $f_{m}:[0,1]\to T_{m}$.
@@
    For each $m\geq 0$ and $J\in\mathscr{E}_{m}$, there exists $w\in\mathcal{A}_{n}^{m}$ such that $f_{m}|J$ is a linear bijection onto either $\phi^{\eta}_{w}(e_{b})$ or $\phi^{\eta}_{w}(e_{l})$, and there exists a unique interval $J^{\prime}\in\mathscr{E}_{m}\setminus\{J\}$ such that $f_{m}|J^{\prime}=(f_{m}|J)\circ\zeta_{J^{\prime}}$, where $\zeta_{J^{\prime}}:J^{\prime}\to J$ is the unique linear orientation reversing map. Furthermore, $f_{m+1}(J)\subset\phi^{\eta}_{w}([0,1)^{2})$. Conversely, if $(p,q)$ is an edge of $T_{m}$, then there exist exactly two intervals $J,J^{\prime}\in\mathscr{E}_{m}$ such that $f_{m}(J)=f_{m}(J^{\prime})=(p,q)$.
@@
    There are disjoint nondegenerate closed intervals $I_{i,1},I_{i,2},I_{i,3},I_{i,4}\subset[0,1]$, denoted by $I_{i,j}=[a_{i,j},b_{i,j}]$, $j=1,\dots,4$ equal to the preimages of the four leaves of $T^{i}$, such that
@@
  • If $\tau\in\{\tilde{\tau}_{1},\tilde{\tau}_{2}\}$, then for a vertex $v\in V$ that is not a leaf, choose a nondegenerate component $I_{v}$ of the preimage $\tau^{-1}(\{v\})$, and for each vertex $v\in V$ that is a leaf other than $(1,0)$ or $(0,1)$, let $I_{v}=I_{i,j}$ where $j\in\{1,2,3,4\}$ is chosen to satisfy $\tau(I_{i,j})=v$. We then define $\mathcal{N}(\tau)=\{I_{v}:v\in V\setminus\{(1,0),(0,1)\}\}$.
@@
Intervals in $\mathscr{E}_{m}$. Let $I\in\mathscr{E}_{m}$ and write $I=(a,b)$. By (P3) there exists $I^{\prime}=(a^{\prime},b^{\prime})\in\mathscr{E}_{m}\setminus\{I\}$ such that $f_{m}|I^{\prime}=(f_{m}|I)\circ\zeta_{I^{\prime}}$ where $\zeta_{I^{\prime}}:I^{\prime}\to I$ is the unique linear orientation reversing map and there exists a word $v\in\mathcal{A}_{n}^{m}$ such that $f_{m}(I)=\phi_{v}^{\eta}(e_{b})$ or $f_{m}(I)=\phi_{v}^{\eta}(e_{l})$. If $f_{m}(I)=\phi^{\eta}_{v}(e_{l})$, then, without loss of generality, we may assume that the $y$-coordinate of $f_{m}$ is increasing inside $I$ and decreasing inside $I^{\prime}$. Let $\tilde{\theta}_{I}:I\to(0,s_{\eta(v),U})$ be linear, bijective, and increasing, and let $\tilde{\xi}_{I^{\prime}}:I^{\prime}\to(s_{\eta(v),U},\tilde{t}_{\eta(v)})$ be linear, bijective, and decreasing. We define
@@
If $J\in\mathscr{N}_{m+1}(I)$, then there is a character $j\in\mathcal{A}_{n}$ such that $f_{m+1}(J)=\phi_{vj}^{\eta}({\bf 0})$, and we define $\operatorname{{\bf w}}_{m+1}(I)=vj$. Note that if $J,Q\in\mathscr{N}_{m+1}(I)\cup\mathscr{N}_{m+1}(I^{\prime})$ are distinct, then $\operatorname{{\bf w}}_{m+1}(J)\neq\operatorname{{\bf w}}_{m+1}(Q)$ as $\tilde{\theta}_{I}(J)\neq\tilde{\theta}_{I}(Q)$ are distinct intervals in, for example, $\mathcal{N}(\tilde{\tau}_{\eta(v)})$ ($\tilde{\theta}_{I}$ may be $\tilde{\xi}_{I^{\prime}}$ instead).
@@
By definition of $\tau_{\eta(\operatorname{{\bf w}}_{m}(I))}$, for any $Q\in\mathscr{N}_{m+1}(I)$, there is a character $j\in\mathcal{A}_{n}$ satisfying $f_{m+1}(Q)=\{\phi^{\eta}_{\operatorname{{\bf w}}_{m}(I)j}({\bf 0})\}$, and we set $\operatorname{{\bf w}}_{m+1}(Q)=\operatorname{{\bf w}}_{m}(I)j$.
@@
To finish the proof, we show that $\operatorname{{\bf w}}_{m+1}:\mathscr{N}_{m+1}\to\mathcal{A}_{n}^{m+1}$ is surjective. Let $u\in\mathcal{A}_{n}^{m+1}$ be a word. By (P1) at $m$, there exists a unique interval $I\in\mathscr{N}_{m}$ such that $\operatorname{{\bf w}}_{m}(I)=u(m)$. If $f_{m}(I)$ is a leaf in $T_{m}$, then by definition of $\mathcal{N}(\tau_{i})$, for each $j\in\mathcal{A}_{n}$ there is an interval $J\in\mathscr{N}_{m+1}(I)$ satisfying $\operatorname{{\bf w}}_{m+1}(J)=\operatorname{{\bf w}}_{m+1}(I)j$, in particular one of them has $\operatorname{{\bf w}}_{m+1}(J)=u$. If $f_{m}(I)$ is not a leaf in $T_{m}$, then the result follows by the definitions of $\mathcal{N}(\tilde{\tau}_{i}),\mathcal{N}(\tau_{i})$ as well as (P3) at $m$. ∎
@@
It is possible that $\mathscr{N}_{m+1}\cap\mathscr{N}_{m}\neq\emptyset$ for some $m\in\mathbb{N}$. Define a map $\operatorname{{\bf w}}:\bigcup_{m\geq 0}\mathscr{N}_{m}\to\mathcal{A}_{n}^{*}$ as follows. Set $\operatorname{{\bf w}}([0,1])=\varepsilon$. If $m\in\mathbb{N}$ and $I\in\mathscr{N}_{m}\setminus\mathscr{N}_{m-1}$, set $\operatorname{{\bf w}}(I)=\operatorname{{\bf w}}_{m}(I)$. If $m\in\mathbb{N}$ and $I\in\mathscr{N}_{m}\cap\mathscr{N}_{m-1}$, set $\operatorname{{\bf w}}(I)=\operatorname{{\bf w}}_{m-k}(I)$, where $k\in\mathbb{N}$ is the greatest integer such that $I\in\mathscr{N}_{m-k}$. Note that $\operatorname{{\bf w}}$ is injective and $V_{m}=\{\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0}):I\in\mathscr{N}_{m}\}$ for every $m\in\mathbb{N}$.
@@
Fix $m\in\mathbb{N}$ and $x\in[0,1]$. By (P1), there exists $I\in\mathscr{N}_{m}\cup\mathscr{E}_{m}\cup\mathscr{F}_{m}$ such that $x\in I$. If $I\in\mathscr{F}_{m}$, then by (P5) we have that $f_{m}(x)=f_{m+1}(x)$. If $I\in\mathscr{N}_{m}$, then by (P4) we have that $f_{m}(x),f_{m+1}(x)\in\phi^{\eta}_{\operatorname{{\bf w}}_{m}(I)}([0,1]^{2})$. Finally, if $I\in\mathscr{E}_{m}$, then by (P3) we have that there is some word $v\in\mathcal{A}_{n}^{m}$ such that $f_{m}(x),f_{m+1}(x)\in\phi^{\eta}_{v}([0,1]^{2})$. In any case, we have the claimed estimate.
Hence, the functions $f_{m}$ converge uniformly to a continuous map $f:[0,1]\to[0,1]^{2}$. Furthermore, because the dendrites $T_{m}\subset K^{\eta}$ are nested and $f_{m}:[0,1]\to T_{m}$ are surjective, we have that the uniform limit $f$ is a continuous parameterization of $K^{\eta}$, as desired. ∎
@@
To prove porosity, note first that by (P3), (P4), and (P6) there exists a word $v\in\mathcal{A}_{n}^{m}$ such that $f(I)\cap f(J)\subset K^{\eta}_{v}$. Fix $x\in K^{\eta}$ and $r\in(0,n^{-m}]$. There exists an integer $k_{0}\in\mathbb{N}$ such that $n^{-k_{0}}\leq r<n^{-k_{0}+1}$, thus there exists $w\in\mathcal{A}_{n}^{k_{0}+1}$ with $K^{\eta}_{w}\subset\overline{B}(x,r)\cap K^{\eta}$.
@@
Therefore, we may assume for the rest of Case 2 that $u=v$ and that neither of $f(I),f(J)$ is a singleton. If $f_{k_{0}+1}(J)\cap K^{\eta}_{w}=\emptyset$, then $f(J)\cap\phi^{\eta}_{w}((0,1)^{2})=\emptyset$ by Lemma 4.2 and by (P3), (P4), and (P5), hence the existence of such characters $i,j\in\mathcal{A}_{n}$ is clear. Similarly if $f_{k_{0}+1}(I)\cap K^{\eta}_{w}=\emptyset$. Now assume that $f_{k_{0}+1}(I)\cap K^{\eta}_{w}\neq\emptyset$ and $f_{k_{0}+1}(J)\cap K^{\eta}_{w}\neq\emptyset$. Then there are distinct leaves $p_{1},p_{2}\in T_{k_{0}+2}$ with $p_{1},p_{2}\in K^{\eta}_{w}$ and distinct intervals $Q_{1},Q_{2}\in\mathscr{N}_{k_{0}+2}$ such that $Q_{1}\subset J$ and $Q_{2}\subset I$ satisfy $f_{k_{0}+2}|Q_{1}\equiv p_{1}$ and $f_{k_{0}+2}|Q_{2}\equiv p_{2}$. Then $f([0,1]\setminus Q_{1})\cap\phi^{\eta}_{\operatorname{{\bf w}}_{k_{0}+2}(Q_{1})}((0,1)^{2})=\emptyset$ and $\operatorname{{\bf w}}(Q_{1})(k_{0}+1)=w$, therefore there is a pair of characters $i,j\in\mathcal{A}_{n}$ such that $wj=\operatorname{{\bf w}}_{k_{0}+2}(Q_{1})$ and $K^{\eta}_{wji}\subset\phi^{\eta}_{wj}((0,1)^{2})$. Since $I,J$ are distinct, and thus disjoint, we have that $f(I)\cap\phi^{\eta}_{wj}((0,1)^{2})=\emptyset$, concluding Case 2.
@@
Note that by (P7), there is some $I\in\mathscr{N}_{m+1}$ such that $f_{m+1}(I)$ is a leaf in $T_{m+1}$ and $I\subset J$. Then since $f_{m+1}(I)=\phi^{\eta}_{\operatorname{{\bf w}}(I)}({\bf 0})$ is a leaf in $T_{m+1}$, we have that
@@
If $Q\in\mathscr{F}_{m+1}$, then by (P5) there exists a word $u\in\mathcal{A}_{n}^{m+1}$ satisfying
@@
If $Q\in\mathscr{E}_{m+1}$, then by (P3) and Lemma 6.4, there exists a word $u\in\mathcal{A}_{n}^{m+1}$ such that either $f_{m+1}(Q)=\phi_{u}^{\eta}(e_{b})$ or $f_{m+1}(Q)=\phi_{u}^{\eta}(e_{l})$. In either case, by (P3), (P4), and (P5) we have that $f_{k}(Q)\subset\phi_{u}^{\eta}([0,1)^{2})$. Note that $\phi_{\operatorname{{\bf w}}(I)}^{\eta}(e_{l})$ emanates from $\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0})$ upward and $\phi_{\operatorname{{\bf w}}(I)}^{\eta}(e_{b})$ emanates from $\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0})$ to the right. Since $\phi_{\operatorname{{\bf w}}(I)}^{\eta}({\bf 0})$ is a leaf in $T_{m+1}$, by Lemma 6.3 we must have that $u\neq\operatorname{{\bf w}}(I)$.
@@
so $\phi^{\eta}_{\operatorname{{\bf w}}(I)}((0,1)^{2})\cap K^{\eta}\subset f(I)$. Since $I$ is compact, we have $K^{\eta}_{\operatorname{{\bf w}}(I)}\subset f(I)$, hence $K^{\eta}_{\operatorname{{\bf w}}(I)}\subset f(J)$ and $|\operatorname{{\bf w}}(I)|=m+1$. Therefore, $\pi_{\eta}\#\nu_{n}(f(J))\geq(5n-6)^{-m-1}$. ∎
@@
Let $f_{m}$, $\mathscr{E}_{m},\mathscr{N}_{m},\mathscr{F}_{m}$, and $\operatorname{{\bf w}}$ be as in Lemma 6.5 and Remark 6.6. There exists a non-decreasing surjection $\zeta:[0,1]\to[0,1]$ such that